OpenStudy (tapplewhaite):

Suppose that a customer service center claims that the telephone calls is receives last 112 seconds on average, with a standard deviation of 9 seconds. If you took a sample of 324 telephone calls, which of the following mean times would be within the 95% confidence interval? A. 115.5 seconds B. 112.5 seconds C. 114.5 seconds D. 113.5 seconds

OpenStudy (holsteremission):

Assuming the duration of calls is normally distributed, the confidence interval takes the form \(\bar x\pm Z_{\alpha/2}\dfrac{\sigma}{\sqrt n}\), where \(\bar x\) is the mean, \(\sigma\) is the standard deviation, \(n\) is the sample size, and \(Z_{\alpha/2}\) is the critical value at \((1-\alpha)\times100\%\) confidence. In other words, \(\mathbb P(|Z|<Z_{\alpha/2})=95\%\), which mean \(Z_{\alpha/2}\approx1.96\). So the confidence interval is \((111.02,112.98)\).